# Part-aware Shape Generation with Latent 3D Diffusion of Neural Voxel Fields

### v3

@@ IV-C Ablation Studies
with PASD 27.70 1.03 60.6 36.01 0.18 82.1
without PASD 29.18 1.05 59.5 38.23 0.21 79.3
Importance of the latent 3D diffusion. Different from previous methods, this paper proposes the latent 3D diffusion that transfers the diffusion process into latent space, increasing the generative resolution significantly and preserving the high-quality generation. We further apply the diffusion process to the voxel field directly to observe the performance change. Due to the limitation of computational cost, we set the resolution of voxel fields to 32$\times$32$\times$32. Tab. IV reports the performance comparison. Without the latent diffusion, the FID drops a lot, and the geometry metric of MMD and COV drop 0.36 and 4.9, which shows the importance of latent diffusion. Moreover, we visualize the qualitative results of the two different settings in Fig. 5, and the results reveal that the latent 3D diffusion brings high-quality geometry and rendering details.
Importance of the part-aware shape decoder. The part-aware shape decoder incorporates the part code into the neural voxel field, generating the structural shape with accurate rendering results. To demonstrate its importance, we replace it with two independent MLPs, rendering the color and part maps from the voxel field directly. Tab. V shows that the part-aware shape decoder can bring more improvements in rendering quality. As depicted in Fig. 4, the part-aware shape decoder helps the model learn the accurate structural information, which allows for consistent rendering results of the local region.
Importance of the gradient skip. The gradient skip technique aims to reduce redundant gradient calculations and computational costs by jointly learning the latent 3D diffusion model and the part-aware shape decoder. In Fig. 11, we compare our method with the split training strategy, which does not utilize the gradient skip technique. It is evident from the comparison that the split training strategy lacking gradient skipping leads to the learning of a biased neural voxel field, consequently yielding inferior rendered images.
@@
Effect of sampling steps. By leveraging rapid sampling from the decoupled diffusion model, we achieve satisfactory results in just 30 steps. Tab. VI presents a comparison of performance based on various sampling steps for chair generation. With an increase in the sampling step, performance shows a gradual improvement. Further increasing the sampling steps results in only a marginal improvement, suggesting that the generation quality is nearing saturation. It is observed that the sampling step has a greater impact on FID and KID metrics, highlighting its influence on image quality rather than geometry.
TABLE VI: Performance comparisons of different sampling steps for generating chairs.
@@ V Conclusion
In this paper, we introduce a method for generating part-aware neural voxel fields using latent 3D diffusion models. The proposed method consists of a latent 3D diffusion model and a part-aware shape decoder. On one hand, the latent 3D diffusion model enhances the high resolution of the neural voxel fields, thereby improving the geometric details and the quality of rendered images. On the other hand, the part-aware shape decoder integrates part codes into voxel fields to generate part-aware shapes with high-quality rendered images. We evaluate the effectiveness of our method on four classes by comparing it against state-of-the-art diffusion-based approaches, demonstrating its efficacy in generating precise part-aware shapes.

### v5

@@ IV-C Ablation Studies
with PASD 27.73 1.00 61.5 35.94 0.18 82.3
without PASD 30.09 1.03 61.0 38.84 0.19 81.8
Importance of the latent 3D diffusion. Different from previous methods, this paper proposes the latent 3D diffusion that transfers the diffusion process into latent space, increasing the generative resolution significantly and preserving the high-quality generation. We further apply the diffusion process to the voxel field directly to observe the performance change. Due to the limitation of computational cost, we set the resolution of voxel fields to 32$\times$32$\times$32. Tab. IV reports the performance comparison. Without the latent diffusion, the FID drops a lot, and the geometry metric of MMD and COV drop 0.40 and 6.4 on Chairs, which shows the importance of latent diffusion. Moreover, we visualize the qualitative results of the two different settings in Fig. 6, and the results reveal that the latent 3D diffusion brings high-quality geometry and rendering details.
Importance of the part-aware shape decoder. The part-aware shape decoder incorporates the part code into the neural voxel field, generating the structural shape with accurate rendering results. To demonstrate its importance, we replace it with two independent MLPs, rendering the color and part maps from the voxel field directly. Tab. V shows that the part-aware shape decoder can bring more improvements in rendering quality (FID) rather than geometry quality (MMD and COV). Since the part-aware shape decoder helps distinguish different local regions instead of the whole object geometry, the MMD and COV metrics improve marginally. As depicted in Fig. 5, the part-aware shape decoder helps the model learn the accurate structural information, which allows for consistent rendering results of the local region. We also plot the visualization of the generated fields in Fig. 11, including the generated fields by the diffusion model and the refined fields by the part-aware shape decoder. The part-aware shape decoder utilizes the cross- and self-attention to build relationships between different voxels, letting the model learn more discriminative features in each local part region and improving the part-aware generation. The red circle in Fig. 11 shows the improvements of local details in different parts and the purple circle shows that the refined fields are smoother.
Figure 11: Comparisons between the voxel fields generated by the diffusion model and refined by the part-aware shape decoder.
TABLE VI: Ablation on different training strategy.
Method Chair Table Airplane Car Guitar Lamp Laptop Pistol
[…]
Separate training 1.24 58.2 1.45 56.9 0.31 79.9 0.70 33.6 0.25 40.5 1.70 56.8 0.89 35.3 0.48 45.5
Joint training 1.00 61.5 1.13 59.9 0.18 82.3 0.46 42.1 0.14 45.1 1.48 59.7 0.79 38.8 0.31 50.5
Importance of the gradient skip. The gradient skip technique aims to reduce redundant gradient calculations and computational costs by jointly learning the latent 3D diffusion model and the part-aware shape decoder. In Fig. 13, we compare our method with the separate training strategy, in which we train the diffusion model first and then the part-aware shape decoder. It is evident from the comparison that the separate training strategy lacking gradient skipping leads to the learning of a biased neural voxel field, consequently yielding inferior rendered images. Tab. VI shows that the separate training strategy gets worse results in all categories.
@@
Effect of different diffusion models and sampling steps. Firstly, we utilize typical diffusion models DDPM [21] and DDIM [51] as our base diffusion model. As shown in Tab. VII, by leveraging rapid sampling from the decoupled diffusion model, our method only takes 10 steps to generate similar quality to DDIM with 30 steps, showing superior efficiency in generating high-quality shapes. Tab. VII presents a comparison of performance based on various sampling steps for chair generation. With an increase in the sampling step, performance shows a gradual improvement. Further increasing the sampling steps results in only a marginal improvement, suggesting that the generation quality is nearing saturation.
TABLE VII: Ablation on different diffusion methods and sampling steps. Note that we report MMD / COV metrics on the Chair category.
@@ V Conclusion
In this paper, we introduce a method for generating part-aware neural voxel fields using latent 3D diffusion models. The proposed method consists of a latent 3D diffusion model and a part-aware shape decoder. On the one hand, the latent 3D diffusion model enhances the high resolution of the neural voxel fields, thereby improving the geometric details and the quality of rendered images. We incorporate part codes into the latent 3D diffusion model, allowing the model to learn part-aware information in field generation. On the other hand, the part-aware shape decoder integrates part codes into voxel fields to generate part-aware shapes with high-quality rendered images. We evaluate the effectiveness of our method on four classes by comparing it against state-of-the-art diffusion-based approaches, demonstrating its efficacy in generating precise part-aware shapes.
VI Acknowledgment
This work is supported in part by the NSFC (62325211, 62132021, 62372457), the Major Program of Xiangjiang Laboratory (23XJ01009), Young Elite Scientists Sponsorship Program by CAST (2023QNRC001), the Natural Science Foundation of Hunan Province of China (2022RC1104) and the NUDT Research Grants (ZK22-52).